# Shafarevich’s conjecture for families of hypersurfaces over function fields

### v1

@@ Abstract.
Given a smooth quasi-projective complex algebraic variety $\mathcal{S}$, we prove that there are only finitely many Hodge-generic non-isotrivial families of smooth projective hypersurfaces over $\mathcal{S}$ of degree $d$ in $\mathbb{P}_{{\mathbb{C}}}^{n+1}$. We prove that the finiteness is uniform in $\mathcal{S}$ and we give examples where the result is sharp. We also prove similar results for certain complete intersections in $\mathbb{P}_{{\mathbb{C}}}^{n+1}$ of higher codimension and more generally for algebraic varieties whose moduli space admits a period map that satisfies the infinitesimal Torelli theorem.
@@ 1. Introduction
Fix an integer $g\geq 1$, a complex projective curve ${\mathcal{S}}$, and a finite set ${\mathcal{T}}$ of points in ${\mathcal{S}}$. Are there only finitely many nonisotrivial families of smooth projective genus $g$ curves over ${\mathcal{S}}\backslash{\mathcal{T}}$?
@@
1.2 was resolved by Parshin and Arakelov [Par68, Ara72], and 1.3 was answered by Faltings for families of abelian varieties over complex curves [Fal83b]. It has also been studied in other situations in [Pet86, Sai93, BV00, LTYZ05, LTYZ11, Kov02, KL10] and more recently in [AT24]. We refer to the introduction of [KL10] for more references and detailed historical background.
@@
See Definition 2.1 for a definition of Hodge-generic, which in this case we take to be with respect to the moduli space of hypersurfaces of degree $d$ and dimension $n$.
@@
Let $\mathcal{S}$ be a smooth quasi-projective curve over ${\mathbb{C}}$. There are finitely many smooth projective non-isotrivial Hodge-generic families of $n$-dimensional hypersurfaces of arbitrary degree $d$ and bounded discriminant over $\mathcal{S}$.
We remark that the assumption on the family being Hodge-generic is mild in light of recent results of [BKU21]. Indeed, the authors proved that for $n\geq 3,d\geq 5$, $(n,d)\neq(4,5)$, there exists a Zariski open subset $U$ in the moduli space of hypersurfaces $\mathcal{U}_{n,d}$ such that any variety ${\mathcal{S}}$ intersecting $U$ nontrivially [^3] I.e., the intersection is nonempty. is Hodge-generic. On the other hand, Saito–Zucker [SZ01, Section 8] and Viehweg–Zuo [VZ05, Theorem 0.3] have given examples of non-Hodge generic families of K3 surfaces [^4] Which can be chosen to be quartic surfaces. and hypersurfaces with $d\geq n+1$ which deform non-trivially in positive dimensional families. Hence the assumption on the Mumford–Tate group in Theorem 1.4 is necessary in general.
@@ 1.1. Main contribution and earlier work
The main observation of this paper is to recast the two problems in the framework of Hodge theory: the boundedness problem is solved using the recent resolution of the conjecture of Griffiths by Brunebarbe-Bakker-Tsimerman [BBT23], and we give a sufficient condition for rigidity in Hodge-theoretic terms and in situations where infinitesimal Torelli holds, directly inspired from earlier results of Peters, see [Pet86, Pet90].
Our main theorem is the following:
@@
Let $\mathcal{S}$ be smooth quasiprojective algebraic variety and $\mathcal{M}$ be a separated, finite type, smooth Deligne–Mumford stack which is a fine moduli space for polarized smooth projective algebraic varieties. Assume that ${\mathcal{M}}$ admits a period map that satisfies the infinitesimal Torelli theorem and such that the associated generic Mumford–Tate group is absolutely simple.
Then Shafarevich finiteness holds for Hodge generic non-isotrivial families over $\mathcal{S}$ parameterized by $\mathcal{M}$. Moreover, the number of such families is bounded uniformly when $\mathcal{S}$ is the fiber of a smooth family ${\mathcal{C}}\rightarrow{\mathcal{T}}$.
@@ 1.2. Strategy of the proof
As explained in Section 1.1, the proof of Theorem 1.4 has the same two ingredients: we first prove that the space of maps $\mathrm{Hom}(\mathcal{S},\mathcal{M})$ has finitely many connected components over ${\mathbb{C}}$. We first show that ${\mathcal{M}}$ is weakly bounded and compactifiable using the generalized Arakelov inequality recently proved by Brunebarbe [Bru20] and the ampleness of the Griffiths line bundle on the coarse moduli space associated to $\mathcal{M}$. Then we invoke Kovács–Lieblich’s result to show that $\mathrm{Hom}(\mathcal{S},\mathcal{M})$ is a finite type stack. As for the rigidity, we use a result of Peters to analyze deformation of period maps and relate that to deformation of maps to $\mathcal{M}$. Under the assumption on the generic Mumford–Tate group being absolutely simple, we prove that these maps are rigid and thus $\mathrm{Hom}(\mathcal{S},\mathcal{M})$ is zero dimensional at those points.
@@ 1.4. Acknowledgments
We thank Hélène Esnault for her interest in our work and for helpful suggestions, and Aaron Landesman for helpful comments. P.E.   was supported by DMS-2201221. S.T. was supported by NSF grant DMS-2302388. A.L. was supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE 2140743. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
@@ 2.1. Hodge-theoretic background
Let $H_{s}$ be the algebraic monodromy group of ${\mathbb{V}}$ at $s$, defined as the connected component of the Zariski closure over ${\mathbb{Q}}$ of the image of the monodromy representation corresponding to the local system ${\mathbb{V}}_{\mathbb{Z}}$:
@@
By a theorem of Deligne and André [Del71, And92], $H_{s}$ is a normal subgroup of $G^{\rm der}$, the derived group of $G$ for very general $s\in{\mathcal{S}}({\mathbb{C}})$.
@@ 2.2. Arakelov inequality
In this section, we recall Arakelov’s inequality, recently established by Brunebarbe [Bru20], which will be crucial to establishing boundedness. For history and background on Arakelov inequalities, see [Vie09, KT24].
@@
Let $\mathcal{S}\hookrightarrow\overline{\mathcal{S}}$ be a smooth compactification with normal crossings and unipotent monodromies around the points in $\overline{\mathcal{S}}\backslash\mathcal{S}$ and let $\overline{L}$ be the Deligne extension to $\overline{\mathcal{S}}$ of the pullback of the Griffiths line bundle $\varphi^{*}L$. Then we have the following crucial inequality originally due to Arakelov [Ara72] in the case of families of curves and generalized for arbitrary non-isotrivial variations of Hodge structure over $\mathcal{S}$, see [Bru20, Corollary 1.10].
@@
where $\chi({\mathcal{S}})$ is the topological Euler characteristic of ${\mathcal{S}}$, the constant $C({\mathbb{V}})$ only depends on the discrete invariants of ${\mathbb{V}}$, that is, the Hodge numbers and their indices.
@@
where $w$ is the length of the Hodge structure ${\mathbb{V}}$.
@@ 2.3.1. Summary of Kovács–Lieblich work
Finally, we say that ${\mathcal{M}}$ is weakly bounded, if we can find $b$, $L$, as above such that ${\mathcal{M}}$ is weakly bounded with respect to $M^{\dagger}$ and $L$ by $b$.
@@ 2.3.2. Deligne–Mumford stacks with immersive period maps
We assume that ${\mathcal{M}}$ supports a polarized variation of Hodge structure ${\mathbb{V}}$ of weight $k$ and for which the period map satisfies infinitesimal Torelli theorem, i.e., the period map is immersive.
@@ 3.1. Rigidity of period maps
be the corresponding period map. Let $T^{h}X_{\Gamma}\subseteq TX_{\Gamma}$ denote the Griffiths distribution. Then Griffiths transversality implies that $\varphi$ determines a point in $\operatorname{Hom}^{||}(\mathcal{S},X_{\Gamma})$, the space of holomorphic maps that are tangent to $T^{h}X_{\Gamma}$. By [Pet90, Corollary 2.6], the set $\operatorname{Hom}^{||}(\mathcal{S},X_{\Gamma})$ has the structure of a finite dimensional analytic variety666with possibly infinitely many irreducible components and its Zariski tangent space at $\varphi$ is described by the following proposition, see [Pet90, Theorem 3.2].
@@
The Zariski tangent space of $\mathrm{Hom}^{||}(\mathcal{S},X_{\Gamma})$ at $\varphi$ is given by the $(-1,1)$ part of the Hodge structure $\mathrm{End}(V_{{\mathbb{C}}})^{\pi_{1}(\mathcal{S},s)}$, where $V_{{\mathbb{C}}}$ is the fiber of ${\mathbb{V}}_{{\mathbb{C}}}$ at a point $s\in{\mathcal{S}}$, and ${\mathbb{V}}$ is the ${\mathbb{Z}}$-local system giving rise to the period map $\varphi$.
@@
Since the monodromy group $H$ is the connected component of the closure of $\pi_{1}({\mathcal{S}})$ in the Zariski topology, we have
$\mathrm{End}(V)^{\pi_{1}(\mathcal{S})}=\mathrm{End}(V)^{\overline{\pi_{1}(\mathcal{S})}}\subseteq\mathrm{End}(V)^{H}\leavevmode\nobreak\ .$
@@
Keep the same notation as above. If $H$ acts absolutely irreducibly on $V$, then the period map is horizontally rigid.
@@
By Proposition 3.1, the Zariski tangent space of horizontal deformations of $\varphi$ is equal to the $(-1,1)$ part of $\mathrm{End}(V_{{\mathbb{C}}})^{\pi_{1}(\mathcal{S})}$. Since $H$ acts absolutely irreducibly on $V$, by Schur’s lemma, $\mathrm{End}(V_{{\mathbb{C}}})^{H}={\mathbb{C}}$, hence the $(-1,1)$ part of the subspace $\mathrm{End}(V_{{\mathbb{C}}})^{\pi_{1}(\mathcal{S})}$ is trivial, which concludes the proof. ∎
@@ 3.2. Rigidity of families parameterized by a DM stack
${{T_{[f]}\operatorname{Hom}_{\mathbb{C}}(\mathcal{S},{\mathcal{M}})}}$${{\operatorname{Hom}_{\mathcal{O}_{\mathcal{S}}}(f^{*}\Omega_{{\mathcal{M}}/{\mathbb{C}}},{\mathcal{O}}_{\mathcal{S}})}}$${{T_{[{\varphi}\circ f]}\operatorname{Hom}_{\mathbb{C}}^{||}(\mathcal{S},X_{\Gamma})}}$${{\operatorname{Hom}_{\mathcal{O}_{\mathcal{S}}}(f^{*}{\varphi}^{*}\Omega^{h}_{X_{\Gamma}/{\mathbb{C}}},{\mathcal{O}}_{\mathcal{S}})}}$
where $\Omega^{h}_{X_{\Gamma}/{\mathbb{C}}}$ is the dual to the Griffiths distribution on $X_{\Gamma}$. Assumption 2.6 states that $\varphi$ is immersive, hence $d\varphi:{\varphi}^{*}\Omega_{X_{\Gamma}/{\mathbb{C}}}\to\Omega_{{\mathcal{M}}/\mathbb{C}}$ is surjective, and so we conclude the map (1) is injective. ∎
@@
The connected components of $\mathrm{Hom}(\mathcal{S},\mathcal{M})$ corresponding to families with absolutely irreducible monodromy representation are zero-dimensional.
@@
$T_{f}\operatorname{Hom}(\mathcal{S},\mathcal{M})\hookrightarrow T_{\varphi\circ f}\mathrm{Hom}^{||}(\mathcal{S},X_{\Gamma}),$
By Proposition 3.1, the space $T_{\varphi\circ f}\mathrm{Hom}^{||}(\mathcal{S},\Gamma\backslash D)$ is identified with the $(-1,1)$-part of the Hodge structure $\mathrm{End}(V)^{\pi_{1}({\mathcal{S}})}$. Now we may apply Proposition 3.3 to conclude that
$T_{\varphi\circ f}\mathrm{Hom}^{||}(\mathcal{S},\Gamma\backslash D)=0,$
@@
In the case that the family corresponding to $f:\mathcal{S}\to\mathcal{M}$ is Hodge-generic and the generic Mumford–Tate group $G$ is absolutely simple, the generic Mumford–Tate group will act on a fiber $V={\mathbb{V}}_{s}$ of ${\mathbb{V}}$ absolutely irreducibly. By Deligne–André, $H$ is a normal subgroup of $G^{\operatorname{der}}$. Since $f:{\mathcal{S}}\to{\mathcal{M}}$ is nonconstant and the $\varphi:M\to X_{\Gamma}$ is immersive, the composition ${\mathcal{S}}\to\Gamma\backslash D$ is non-constant. [^7] In fact, the monodromy representation $\pi_{1}({\mathcal{S}},s)\to\operatorname{GL}(V)$ cannot have finite image, otherwise there is a finite cover ${\mathcal{S}}^{\prime}\to{\mathcal{S}}$ such that the pullback local system is trivial, contradicting the nonconstancy of the map ${\mathcal{S}}\to\mathcal{M}$. Therefore the Zariski closure of the $\pi_{1}$ image is nonfinite and hence positive-dimensional. In particular, the monodromy group $H$ is nontrivial, and therefore $H=G^{\mathrm{der}}$, so $H$ also acts absolutely irreducibly on $V$, and we have rigidity of $f$. Combining this rigidity result with the boundedness result of Section 2.3 concludes the proof of the finiteness proof of Theorem 1.8.
@@
Observe that the assumption of Hodge-genericity in the theorem statements serves primarily to ensure that the monodromy representation is absolutely irreducible, which was really the key assumption.

### v2

@@ Abstract.
Given a smooth quasi-projective complex algebraic variety $\mathcal{S}$, we prove that there are only finitely many Hodge-generic non-isotrivial families of smooth projective hypersurfaces over $\mathcal{S}$ of degree $d$ in $\mathbb{P}_{{\mathbb{C}}}^{n+1}$. We prove that the finiteness is uniform in $\mathcal{S}$ and give examples where the result is sharp.
We also prove similar results for certain complete intersections in $\mathbb{P}_{{\mathbb{C}}}^{n+1}$ of higher codimension and more generally for algebraic varieties whose moduli space admits a period map that satisfies the infinitesimal Torelli theorem.
@@ 1. Introduction
Fix an integer $g\geq 1$, a complex projective curve ${\mathcal{S}}$, and a finite set ${\mathcal{T}}$ of points in ${\mathcal{S}}$. Are there only finitely many non-isotrivial families of smooth projective genus $g$ curves over ${\mathcal{S}}\backslash{\mathcal{T}}$?
@@
1.2 was resolved by Parshin and Arakelov [Par68, Ara72], and 1.3 was answered by Faltings for families of abelian varieties over complex curves [Fal83b]. It has also been studied in other situations in [Pet86, Sai93, BV00, LTYZ05, LTYZ11, Kov02, KL10, JL24] and more recently in [AT24, CHSZ24]. We refer to the introduction of [KL10] for more references and detailed historical background. Throughout the paper, the finite set of such families could very well be empty.
@@
We refer to Definition 2.1 for a definition of Hodge-generic, which in this case we take to be with respect to the moduli space of hypersurfaces of degree $d$ and dimension $n$.
@@
Let $\mathcal{S}$ be a smooth quasi-projective curve over ${\mathbb{C}}$. Then there are only finitely many smooth projective non-isotrivial Hodge-generic families of $n$-dimensional hypersurfaces of arbitrary degree $d$ and bounded discriminant over $\mathcal{S}$.
We remark that the assumption on the family being Hodge-generic is mild in light of recent results of [BKU24].
Indeed, the authors proved that for $n\geq 3,d\geq 5$, $(n,d)\neq(4,5)$, there exists a Zariski open subset $U$  [^3] The set $U$ is the complement of certain unspecified maximal components of the Hodge locus. in the moduli space of hypersurfaces $\mathcal{U}_{n,d}$ such that any variety ${\mathcal{S}}$ intersecting $U$ nontrivially [^4] I.e., the intersection is nonempty. is Hodge-generic. On the other hand, Saito–Zucker [SZ01, Section 8] and Viehweg–Zuo [VZ05, Theorem 0.3] have given examples of non-Hodge generic families of K3 surfaces [^5] Which can be chosen to be quartic surfaces. and hypersurfaces with $d\geq n+1$ which deform non-trivially in positive dimensional families. Hence the assumption on the Mumford–Tate group in Theorem 1.4 is necessary in general.
@@
Remark 1.8.
Theorems 1.4 and 1.7 generalize to hypersurfaces in Severi–Brauer schemes over ${\mathcal{S}}$.
@@ 1.1. Main contribution and earlier work
The main observation of this paper is to recast the two problems in the framework of Hodge theory: the boundedness problem is solved using the recent resolution of the conjecture of Griffiths by Brunebarbe-Bakker-Tsimerman [BBT23], and we give a sufficient condition for rigidity in Hodge-theoretic terms and in situations where infinitesimal Torelli holds, directly inspired by earlier results of Peters, see [Pet86, Pet90]. When ${\mathcal{M}}$ is a variety, boundedness has also been previously obtained in [JL24, Theorem 4.2].
Our main theorem is the following.
@@
Let $\mathcal{S}$ be a smooth quasiprojective algebraic variety and $\mathcal{M}$ be a separated, finite type, smooth Deligne–Mumford stack which is a fine moduli space for polarized projective algebraic varieties. Assume that ${\mathcal{M}}$ admits a period map that satisfies the infinitesimal Torelli theorem, and let $H$ be its algebraic monodromy group.
Then Shafarevich finiteness holds for families over $\mathcal{S}$ parameterized by $\mathcal{M}$666i.e., elements of $\mathrm{Hom}({\mathcal{S}},{\mathcal{M}})$. and whose algebraic monodromy group is equal to $H$.
Moreover, for a smooth family ${\mathcal{C}}\rightarrow{\mathcal{T}}$, the number of families over ${\mathcal{C}}_{t}$ for $t\in T$ satisfying the monodromy condition above is uniformly bounded in $t\in T$.
Independently, Chen, Hu, Sun, and Zuo proved similar results in [CHSZ24].
@@ 1.2. Strategy of the proof
As explained in Section 1.1, the proof of Theorem 1.4 has two ingredients: we first prove that the space of maps $\mathrm{Hom}(\mathcal{S},\mathcal{M})$ is of finite type over ${\mathbb{C}}$ using a method inspired from Kovács and Lieblich. We first show that ${\mathcal{M}}$ is weakly bounded and compactifiable using the Arakelov inequality proved by Peters [Pet86], see also Brunebarbe [Bru20] for a generalization, and the ampleness of the Griffiths line bundle on the coarse moduli space associated to $\mathcal{M}$ established recently in [BBT23]. As for the rigidity, we use a result of Peters [Pet90] to analyze deformation of period maps and relate that to deformation of maps to $\mathcal{M}$. When the algebraic monodromy is maximal, i.e., equal to the monodromy of the universal family, we prove that these maps are rigid and thus $\mathrm{Hom}(\mathcal{S},\mathcal{M})$ is zero dimensional at those points.
@@ 1.4. Acknowledgments
We thank Hélène Esnault for her interest in our work and for helpful suggestions, Aaron Landesman for suggesting Remark 1.8, Ariyan Javanpeykar and Chris Peters for helpful comments and suggestions, and Matt Kerr for pointing out an error in Section 3.3. We thank the referee for valuable suggestions. P.E.   was supported by DMS-2201221. S.T. was supported by NSF grant DMS-2302388 and NSF grant DMS-2503815. A.L. was supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE 2140743.
@@ 2.1. Hodge-theoretic background
Let $H$ be the algebraic monodromy group of ${\mathbb{V}}$ at $s$, defined as the connected component of the Zariski closure over ${\mathbb{Q}}$ of the image of the monodromy representation corresponding to the local system ${\mathbb{V}}_{\mathbb{Z}}$:
@@
By a theorem of Deligne and André [Del71, And92], $H$ is a normal subgroup of $G^{\rm der}$, the derived group of $G$ for very general $s\in{\mathcal{S}}({\mathbb{C}})$.
@@ 2.2. Arakelov inequality
In this section, we recall Arakelov’s inequality which will be crucial to establishing boundedness. For history and background on Arakelov inequalities, see [Vie09, KT24].
@@
Let $\mathcal{S}\hookrightarrow\overline{\mathcal{S}}$ be a smooth compactification of ${\mathcal{S}}$ and unipotent monodromies around the points in $\overline{\mathcal{S}}\backslash\mathcal{S}$ and let $\overline{L}$ be the Deligne extension to $\overline{\mathcal{S}}$ of the pullback of the Griffiths line bundle $\varphi^{*}L$. Then we have the following crucial inequality originally due to Arakelov [Ara72] in the case of families of curves and generalized for arbitrary non-isotrivial variations of Hodge structure over $\mathcal{S}$, see [Pet00] and [Bru20, Corollary 1.10]. When ${\mathcal{S}}$ has higher dimension, see [Bru20].
@@
where $\chi({\mathcal{S}})$ is the topological Euler characteristic of ${\mathcal{S}}$ and the constant $C({\mathbb{V}})$ only depends on the discrete invariants of ${\mathbb{V}}$, that is, the Hodge numbers and their indices.
@@
where $w$ is the level  [^7] By definition, if $[a,b]$ is the smallest interval such that $h^{i,n-i}=0$ for $i\notin[a,b]$, then the level of ${\mathbb{V}}$ is the integer $b-a$. of the Hodge structure ${\mathbb{V}}$.
@@ 2.3.1. Summary of Kovács–Lieblich work
Finally, we say that ${\mathcal{M}}$ is weakly bounded, if we can find $b$ and $L$ as above such that ${\mathcal{M}}$ is weakly bounded with respect to $M^{\dagger}$ and $L$ by $b$.
@@ 2.3.2. Deligne–Mumford stacks with immersive period maps
We assume that ${\mathcal{M}}$ supports a polarized variation of Hodge structure ${\mathbb{V}}$ of weight $k$ and for which the period map satisfies an infinitesimal Torelli theorem, i.e. the period map is immersive.
@@ 3.1. Rigidity of period maps
be the corresponding period map. Let $T^{h}X_{\Gamma}\subseteq TX_{\Gamma}$ denote the Griffiths distribution. Then Griffiths transversality implies that $\varphi$ determines a point in $\operatorname{Hom}^{||}(\mathcal{S},X_{\Gamma})$, the space of holomorphic maps that are tangent to $T^{h}X_{\Gamma}$. By [Pet90, Corollary 2.6], the set $\operatorname{Hom}^{||}(\mathcal{S},X_{\Gamma})$ has the structure of a finite-dimensional analytic variety999with possibly infinitely many irreducible components and its Zariski tangent space at $\varphi$ is described by the following proposition, see [KP16, (10.6)]. [^10] See also [Pet90, Theorem 3.2] for earlier work.
@@
The Zariski tangent space of $\mathrm{Hom}^{||}(\mathcal{S},X_{\Gamma})$ at $\varphi$ is given by the $(-1,1)$ part of the Hodge structure $\mathfrak{g}_{\mathbb{C}}^{\pi_{1}(\mathcal{S},s)}$, where $\mathfrak{g}\subset\mathrm{End}(V)$ is the Lie algebra of $G$.
@@
Since the geometric monodromy group $H$ is the connected component of the closure of $\rho(\pi_{1}({\mathcal{S}}))$ in the Zariski topology, we have
$\mathfrak{g}^{\pi_{1}(\mathcal{S})}=\mathfrak{g}^{\overline{\pi_{1}(\mathcal{S})}}\subseteq\mathfrak{g}^{H}\leavevmode\nobreak\ .$
@@
Keep the same notation as above. If $H=G^{\operatorname{der}}$, the derived group of $G$, then the period map of ${\mathcal{S}}$ is horizontally rigid.
@@
By Proposition 3.1, the Zariski tangent space of horizontal deformations of $\varphi$ is equal to the $(-1,1)$ part of $\mathfrak{g}^{\pi_{1}(\mathcal{S})}$. Since $G$ is reductive, the invariant part $\mathfrak{g}^{\pi_{1}(\mathcal{S})}$ is purely of type $(0,0)$, equal to the Lie algebra of its center, which is Hodge–Tate. Therefore, the $(-1,1)$ part is trivial. ∎
@@ 3.2. Rigidity of families parameterized by a DM stack
${{T_{f}\operatorname{Hom}_{\mathbb{C}}(\mathcal{S},{\mathcal{M}})}}$${{\operatorname{Hom}_{\mathcal{O}_{\mathcal{S}}}(f^{*}\Omega_{{\mathcal{M}}/{\mathbb{C}}},{\mathcal{O}}_{\mathcal{S}})}}$${{T_{{\varphi}\circ f}\operatorname{Hom}_{\mathbb{C}}^{||}(\mathcal{S},X_{\Gamma})}}$${{\operatorname{Hom}_{\mathcal{O}_{\mathcal{S}}}(f^{*}{\varphi}^{*}\Omega^{h}_{X_{\Gamma}/{\mathbb{C}}},{\mathcal{O}}_{\mathcal{S}})}}$
where $\Omega^{h}_{X_{\Gamma}/{\mathbb{C}}}$ is the quotient by the annihilator of the horizontal distribution on $X_{\Gamma}$. Assumption 2.6 states that $\varphi$ is immersive, hence $d\varphi:{\varphi}^{*}\Omega_{X_{\Gamma}/{\mathbb{C}}}\to\Omega_{{\mathcal{M}}/\mathbb{C}}$ is surjective, and so we conclude the map (1) is injective. ∎
@@
The connected components of $\mathrm{Hom}(\mathcal{S},\mathcal{M})$ corresponding to families with monodromy group equal to $G^{\operatorname{der}}$ are zero-dimensional.
@@
$T_{f}\operatorname{Hom}(\mathcal{S},\mathcal{M})\hookrightarrow T_{\varphi\circ f}\mathrm{Hom}^{||}(\mathcal{S},X_{\Gamma}).$
By Proposition 3.1, the space $T_{\varphi\circ f}\mathrm{Hom}^{||}(\mathcal{S},\Gamma\backslash{\mathcal{D}})$ is identified with the $(-1,1)$-part of the Hodge structure $\mathfrak{g}_{\mathbb{C}}^{\pi_{1}({\mathcal{S}})}$. Now we may apply Proposition 3.3 to conclude that
$T_{\varphi\circ f}\mathrm{Hom}^{||}(\mathcal{S},\Gamma\backslash{\mathcal{D}})=0,$
@@
Let $G$ be the generic Mumford-Tate group and let $H$ the monodromy group underlying the variation of Hodge structure on ${\mathcal{M}}$. We have a decomposition of adjoint groups over ${\mathbb{Q}}$: $G^{\operatorname{ad}}=H^{\operatorname{ad}}\times H^{\operatorname{fix}}$, which induces a decomposition of period domains ${\mathcal{D}}={\mathcal{D}}_{H}\times{\mathcal{D}}_{fix}$. Moreover, the period map also decomposes, up to a finite étale cover of ${\mathcal{M}}$, as:
$\Phi:{\mathcal{M}}\rightarrow\Gamma_{H}\backslash{\mathcal{D}}_{H}\times\{*\}$
where the first projection still satisfies infinitesimal Torelli and is rigid. Therefore, we can assume without loss of generality that $H=G^{\operatorname{der}}$. Let $f:\mathcal{S}\to\mathcal{M}$ be a moduli map corresponding to a family parametrized by ${\mathcal{S}}$ and such that its monodromy group is equal to $H=G^{\operatorname{der}}$. By Proposition 3.3 and Proposition 3.4, the map $f$ is rigid. Combining this rigidity result with the boundedness result of Section 2.3 concludes the proof of the finiteness proof of Theorem 1.9.
@@
If the generic Mumford–Tate group $G$ is simple, then any non-isotrivial Hodge generic family will automatically have maximal monodromy, by Deligne–André’s theorem. More generally, rigidity holds whenever the algebraic monodromy of ${\mathcal{S}}$ acts absolutely irreducibly on the Lie algebra of the monodromy of the universal family.